PLEASE HELP SOLVE
FACTOR EACH POLYNOMIAL
Z^3-125P^3

Answer by venugopalramana(3286)

PLEASE HELP SOLVE
FACTOR EACH POLYNOMIAL
Z^3-125P^3
use formula
a^3-b^3=(a-b)(a^2+ab+b^2)
(z)^3-(5p)^3=[z-5p][z^2+5zp+25p^2]